OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

OpenStudy (anonymous):

8x-12=4x+24 8x-4x=24+12 4x=36 x=9

OpenStudy (anonymous):

8x-4x=4x 24+12=36 4x=36 4/36=9 x=9